OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 3x – y = 7 2x + 5y = 4

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

17 39 28 -2

OpenStudy (anonymous):

@bman123

OpenStudy (anonymous):

what do you mean by y [dy] for the set of eq. ??

OpenStudy (anonymous):

hey @FlyinSolo_424

OpenStudy (anonymous):

oh sorry. and like d (little y) instead of ^y it would be the other way

OpenStudy (anonymous):



OpenStudy (anonymous):

@exploringphysics

OpenStudy (anonymous):

list the coefficients without the \(y\) you get \[3,7\]\[2,4\] then \(D_y=3\times 4-7\times 2\) if i remember correctly

OpenStudy (anonymous):

im confused @satellite73

OpenStudy (anonymous):

did you see how i got this \[3,7\]\[2,4\]?

OpenStudy (anonymous):

[3 7] [2 4] id required one can you solve this ?

OpenStudy (anonymous):

We start out with the system:\[ \begin{array}{cccc} 3x &–& y &=& 7 \\ 2x &+& 5y &=& 4 \end{array} \] We have our matrix by taking coefficients: \[ \begin{bmatrix} 3&-1\\ 2&5 \end{bmatrix} \]We have our answer column: \[ \begin{bmatrix} 7\\4 \end{bmatrix} \]We replace the y column with the answer column:\[ \begin{bmatrix} 3&7\\ 2&4 \end{bmatrix} \]Then we find the determinant:\[ 3\times 4 - 7 \times 2 \]